OpenStudy (anonymous):

(4x+1)+(x^2-5) = x^2+4x-4, right?

OpenStudy (anteater):

Yes :)

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anteater):

You're welcome! Have any others you want to check?

OpenStudy (anonymous):

yeah

OpenStudy (anteater):

If you want to type them, I can check them.

OpenStudy (anonymous):

ok i will thanks :P

OpenStudy (anonymous):

(x^2-5)/(4x+1) for (g/f)(x) = (x^2-5)/(4x+1), x=/= -1/4

OpenStudy (anteater):

So, g(x) is x^2 - 5 and f(x) is 4x + 1?

OpenStudy (anonymous):

yeah

OpenStudy (anteater):

And they want you to find (g/f)(x) ? Then yes, that would be your answer. Then, for the second part, are they asking you to find (g/f)(-1/4) ?

OpenStudy (anonymous):

no, just the first part. thanks

OpenStudy (anteater):

Ok :) Got other questions?

OpenStudy (anonymous):

yeah a couple

OpenStudy (anteater):

Ok, I'll wait

OpenStudy (anonymous):

Thanks. For f(x)=4x+1 and g(x)= x^2-5, what is (f x g) (x). Would the answer be 4x^3+x^2-4x-6?

OpenStudy (anteater):

Just a moment, please. :)

OpenStudy (anonymous):

k

OpenStudy (anteater):

Close. If you distribute 4x across x^2 - 5, that gives you the terms 4x^3 - 20x Then if you distribute the 1, that gives you x^2 - 5

OpenStudy (anteater):

So the 4x^3 and x^2 terms are correct

OpenStudy (anonymous):

oh, so it's 4x^3+x^2-20x-5?

OpenStudy (anteater):

Yes :)

OpenStudy (anonymous):

yay thankss

OpenStudy (anteater):

You're welcome!

OpenStudy (anteater):

Got another one?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

for f(x)=2x+1 and g(f) = x^2-7, find (f/g)(x) the answer would be (2x+1)/(x^2-7), x=/= sqrt7 ?

OpenStudy (anonymous):

*x=/= +/- sqrt7

OpenStudy (anteater):

Good!

OpenStudy (anteater):

:D

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

for f(x)=3x+1 and g(x)= x^2-6, find (fog)(4), would i plug in 4 for all x's?

OpenStudy (anteater):

Actually, you can think of these as "nested" functions. (fog)(x) = f(g(x)), so you first find g(4), then take the result and find f of that.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

31?

OpenStudy (anteater):

Let me check :) g(4) = 4^2 - 6 = 16 - 6 = 10, and f(10) = 3(10) + 1 = 31 ... so yes!!!

OpenStudy (anonymous):

woooh!

OpenStudy (anteater):

So output of first function g(x), becomes input for f. :)

OpenStudy (anonymous):

thanks :P

OpenStudy (anonymous):

(3x+1)-(x^2-6) = -x^2+3x-5?

OpenStudy (anteater):

When you distribute the "-" (or -1) across the terms in the second set of parentheses, you have -x^2 + 6 ;)

OpenStudy (anonymous):

so it's x^2-3x-7??

OpenStudy (anteater):

Close :) -x^2 +3x + 7

OpenStudy (anteater):

This might clarify it: (3x + 1) - (x^2 - 6) = 3x + 1 -x^2 + 6 = -x^2 + 3x + 7

OpenStudy (anonymous):

ohh, i get it. could it be written as 7+3x-x^2?

OpenStudy (anteater):

Yes, certainly. Although generally speaking, polynomials are written with the terms in descending order. So the term with the variable which has the largest exponent would be listed first, and the constant term would be last.

OpenStudy (anteater):

But 7 + 3x - x^2 is certainly an equivalent expression.

OpenStudy (anonymous):

i see, thank you :P

OpenStudy (anonymous):

i have 2 more, is that ok?

OpenStudy (anteater):

Wlcm :)

OpenStudy (anteater):

Yes

OpenStudy (anonymous):

if h(x)=(fog)(x) and h(x)= 3root x+3, find g(x) if f(x)= 3root x+2. i'm completely lost on this one

OpenStudy (anteater):

Please tell me is the "+3" under the radical for h(x)?

OpenStudy (anteater):

Or is it like this:\[\sqrt{3x} + 3\]

OpenStudy (anonymous):

it's under the radical

OpenStudy (anteater):

And for f(x), is the 2 under the radical?

OpenStudy (anonymous):

yeah

OpenStudy (anteater):

So, h(x) = \[\sqrt{3x + 3}\] , and f(x) = \[\sqrt{3x + 2}\]

OpenStudy (anonymous):

yeah

OpenStudy (anteater):

Ok, thanks :)

OpenStudy (anteater):

Thinking for a moment.

OpenStudy (anonymous):

k :P

OpenStudy (anteater):

Looking at the expressions for h(x) and f(x), what is different between the two?

OpenStudy (anonymous):

+2 and +3

OpenStudy (anteater):

Right! So, basically what you want to do is find an algebraic expression you could plug in for x, in 3x + 2, so that you would get 3x + 3. I can't see a good way of explaining this other than to give the answer and show you, if that is ok?

OpenStudy (anonymous):

sure

OpenStudy (anteater):

If we let g(x) be x + 1/3 , then if we put in x + 1/3 for our input in f(x), we would have: \[\sqrt{3(x+1/3) + 2}\]

OpenStudy (anteater):

Then, when you distribute the 3 and simplify the radicand, what will you have?

OpenStudy (anonymous):

3x+3?

OpenStudy (anteater):

Yes :)

OpenStudy (anteater):

So, basically what I was looking for was what expression could we put in for x in f(x), so that we would end up with 3x + 3 under the radical. Does that make sense?

OpenStudy (anonymous):

and there would be no radicand left?

OpenStudy (anteater):

No, the square root sign would still be there, but yes, you would have 3x + 3 under the radical.

OpenStudy (anonymous):

yeah, that makes sense

OpenStudy (anonymous):

oh, ok

OpenStudy (anteater):

So we could say g(x) = x + 1/3

OpenStudy (anonymous):

i understand :P

OpenStudy (anteater):

Then f(g(x)) = \[\sqrt{3(x + 1/3) + 2}\] = \[\sqrt{3x + 3} = h(x)\]

OpenStudy (anonymous):

yeah i just got that, thanks =]

OpenStudy (anteater):

You're welcome! Any others?

OpenStudy (anonymous):

1 more

OpenStudy (anteater):

ok

OpenStudy (anonymous):

if f(x)=3x, g(x)=x-3, and h(x)= rootx, find (fogoh)(25)

OpenStudy (anteater):

Let me know what you get. :)

OpenStudy (anonymous):

would i begin like 3 (root(x-3)) ? but plug in 25 for x?

OpenStudy (anteater):

You can think of it as f(g(h(x))), so if you start with h(25) and go from there.

OpenStudy (anteater):

h(25) = 5 ... then g(5) = ..

OpenStudy (anonymous):

is the answer 6?

OpenStudy (anteater):

Yes :D

OpenStudy (anteater):

But let's look at what you wrote as well, because that is also correct

OpenStudy (anonymous):

Ok

OpenStudy (anteater):

You could start like: 3 (rootx - 3) and that should work as well

OpenStudy (anteater):

since root 25 would be 5, then if you subtract 3 that would give 2, and then finally multiply by 2 to get 6.

OpenStudy (anonymous):

alright, cool! :D

OpenStudy (anteater):

I mean it would be \[3(\sqrt{x} - 3)\]

OpenStudy (anonymous):

ok

OpenStudy (anteater):

Good! :D

OpenStudy (anonymous):

those are all the questions i have! thank you so much, i'm ready for my test now :D

OpenStudy (anteater):

Good :) If you would like a little more practice, there are some here at the bottom of the page: http://www.mathsisfun.com/sets/functions-composition.html